OpenStudy (anonymous):

How would you turn this into an equivalent factored form: x^2-100

OpenStudy (nenadmatematika):

=(x-10)(x+10)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

x^2 - 10^2 = (x - 10) ( x + 10)